OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point m=8, (7,1)

OpenStudy (anonymous):

y=mx+b 1=8(7)+b -55=b Y=8x-55

OpenStudy (anonymous):

y=8x+b through(7.1) so 1=56+b b=-55 so y=8x-55